# Simulation of charged nanotubes self-assembly during evaporation of a sessile droplet on a substrate

### v1

@@ IV Conclusion
The results of the numerical simulations indicate that the combined effect of advection, diffusion, and electrostatics, under certain values of key system parameters ($N_{p}$, $t_{\mathrm{max}}$, and $kq^{2}$) leads to a transition from order to disorder in the morphology of the nanotube ring deposit, associated with the coffee-ring effect, which was previously observed in the experiment [8]. The concentration of particles affects the average distance between nanotubes in a solution. The strength of the electrostatic interaction between the nanotubes depends on the distance between them and the magnitude of their charge. Such interaction can lead to the mutual alignment of nanotubes with respect to each other at the beginning of the process. The evaporation time affects the velocity of capillary flow, which transfers particles and rotates them when they reach the periphery of a droplet, where the thickness of the liquid layer corresponds to the diameter of nanotubes. This boundary is known as the fixing radius, $R_{f}$ [17]. Each nanotube that crosses this boundary is deposited on the substrate (i.e., it comes into contact with the substrate and sticks to it due to adhesion). In the model, the boundary $R_{f}$ moves as the liquid evaporates until it collapses at the central point of the region. In reality, the situation may be more complicated due to the appearance of many local breaks (dry spots) in an ultrathin film. However, this does not affect the results of the numerical simulations presented here, as such local ruptures, if they occur, appear only after the ring-shaped deposit has been formed. For further research, a modified method with restricted evaporation is also interesting, when a droplet is located between two substrates [25]. This experiment showed the possibility of obtaining aligned nanotube precipitation with an area of several square centimeters. In addition, it is necessary to develop 3D models for a more accurate description of these processes.
@@ Appendix A Rotation of the nanotube by the fluid flow
Consider the rotation relative to the center of nanotube mass at point $C$. The moment of the force exerted on a particle by a liquid relative to a given point is equal to
@@
$\frac{dv_{x}}{dt}=\frac{R^{2}x}{4\left(x^{2}+y^{2}\right)\left(t_{max}-t\right)^{2}}\left(\frac{1}{\sqrt{1-\frac{x^{2}+y^{2}}{R^{2}}}}-1+\frac{x^{2}+y^{2}}{R^{2}}\right),$ (13)
$\frac{dv_{y}}{dt}=\frac{R^{2}y}{4\left(x^{2}+y^{2}\right)\left(t_{max}-t\right)^{2}}\left(\frac{1}{\sqrt{1-\frac{x^{2}+y^{2}}{R^{2}}}}-1+\frac{x^{2}+y^{2}}{R^{2}}\right).$ (14)
@@
The position of the particle is determined by the coordinates of its center of mass $(x_{C},\,y_{C})$ and the angle $\alpha$ between the nanotube and the $OX$ axis. Let us make the following substitutions: $y_{A}=\tan(\alpha)\cdot x_{A}+b,\quad dl=dx_{A}/\cos(\alpha),\quad\text{where}\quad b=y_{C}-\tan(\alpha)\cdot x_{C}$, and compute the integral,
$\vec{M}=\vec{n}_{z}\frac{\frac{m}{l}R^{2}}{4\left(t_{max}-t\right)^{2}}\cdot\\ \int_{x_{A_{1}}}^{x_{A_{2}}}\frac{b\left(x_{A}-x_{C}\right)}{\left(\frac{x_{A}}{\cos(\alpha)}+b\sin(\alpha)\right)^{2}+b^{2}\cos^{2}(\alpha)}\cdot\left(\frac{1}{\sqrt{1-\frac{\left(\frac{x_{A}}{\cos(\alpha)}+b\sin(\alpha)\right)^{2}+b^{2}\cos^{2}(\alpha)}{R^{2}}}}-1\right.+\\ +\left.\frac{\left(\frac{x_{A}}{\cos(\alpha)}+b\sin(\alpha)\right)^{2}+b^{2}\cos^{2}(\alpha)}{R^{2}}\right)\frac{dx_{A}}{\cos(\alpha)}=\vec{n}_{z}\frac{\frac{m}{l}R^{2}b}{4\left(t_{max}-t\right)^{2}}\cdot\\ \left(\cos(\alpha)\left(-2\operatorname{arctanh}\sqrt{1-g_{A2}}+g_{A2}-\ln\left(g_{A2}\right)+2\operatorname{arctanh}\sqrt{1-g_{A1}}-g_{A1}+\ln\left(g_{A1}\right)\right)\right.-\\ -\frac{b\sin(\alpha)\cos(\alpha)+x_{C}}{R}\cdot\left(\frac{1}{\sqrt{W}}\arctan\left(\frac{Wg_{A2}-2W+g_{A2}}{2\sqrt{W\left(1-g_{A2}\right)\left(g_{A2}-W\right)}}\right)\right.+\\ +2\sqrt{g_{A2}-W}-\frac{2}{\sqrt{W}}\arctan\left(\frac{\sqrt{g_{A2}-W}}{\sqrt{W}}\right)-\frac{1}{\sqrt{W}}\arctan\left(\frac{Wg_{A1}-2W+g_{A1}}{2\sqrt{W\left(1-g_{A1}\right)\left(g_{A1}-W\right)}}\right)-\\ -\left.\left.2\sqrt{g_{A1}-W}+\frac{2}{\sqrt{W}}\arctan\left(\frac{\sqrt{g_{A1}-W}}{\sqrt{W}}\right)\right)\right),$ (16)
@@
$g_{A1}=\frac{\left(\frac{x_{A1}}{\cos(\alpha)}+b\sin(\alpha)\right)^{2}+b^{2}\cos^{2}(\alpha)}{R^{2}},\quad g_{A2}=\frac{\left(\frac{x_{A2}}{\cos(\alpha)}+b\sin(\alpha)\right)^{2}+b^{2}\cos^{2}(\alpha)}{R^{2}},$
$W=\left(\frac{b\cos(\alpha)}{R}\right)^{2}.$
@@
$\vec{M}=\int_{A_{1}C}\left[\overrightarrow{CA},\,dm\left.\frac{d\vec{v}}{dt}\right|_{A}\right]+\int_{CA_{2}}\left[\overrightarrow{CA},\,dm\left.\frac{d\vec{v}}{dt}\right|_{A}\right]\approx\left[\overrightarrow{CS_{1}},\,\left.\frac{d\vec{v}}{dt}\right|_{S_{1}}\right]\frac{m}{2}+\left[\overrightarrow{CS_{2}},\,\left.\frac{d\vec{v}}{dt}\right|_{S_{2}}\right]\frac{m}{2}=\\ =\vec{n}_{z}\frac{mR^{2}}{8\left(t_{max}-t\right)^{2}}\cdot\\ \left(\frac{y_{S_{1}}\left(x_{S_{1}}-x_{C}\right)-x_{S_{1}}\left(y_{S_{1}}-y_{C}\right)}{x_{S_{1}}^{2}+y_{S_{1}}^{2}}\left(\frac{1}{\sqrt{1-\frac{x_{S_{1}}^{2}+y_{S_{1}}^{2}}{R^{2}}}}-1+\frac{x_{S_{1}}^{2}+y_{S_{1}}^{2}}{R^{2}}\right)\right.+\\ +\left.\frac{y_{S_{2}}\left(x_{S_{2}}-x_{C}\right)-x_{S_{2}}\left(y_{S_{2}}-y_{C}\right)}{x_{S_{2}}^{2}+y_{S_{2}}^{2}}\left(\frac{1}{\sqrt{1-\frac{x_{S_{2}}^{2}+y_{S_{2}}^{2}}{R^{2}}}}-1+\frac{x_{S_{2}}^{2}+y_{S_{2}}^{2}}{R^{2}}\right)\right),$ (17)
where $S_{1}$ and $S_{2}$ are some points on the nanotube located on opposite sides of its center.
A comparative analysis of the formulas for the torque acting on a particle from the liquid relative to the center of a nanotube is presented in the figures 8, 9 for the time moment $t=100$ s. Note that $t_{max}=200$ s, $M_{1}$ is calculated by the formula (16), $M_{2},\,M_{3}$ are calculated by the formula (17). For points $S_{1}$ and $S_{2}$ in $M_{2}$, the midpoints of the nanotube halves (points $E$ and $G$) have been used; in $M_{3}$, two points at the ends of the nanotube ($A_{1}$ and $A_{2}$) have been used. The torque calculated using the approximate formula (17) exhibits fairly good accuracy and can be employed for calculations in place of the exact formula (16), which slows down the computation process.
Figure 8: Dependence of the torque acting on a particle from the liquid, relative to the center of the nanotube, on the angle $\alpha$ ($x_{C}=10^{-3}$ m, $y_{C}=0.5\cdot 10^{-3}$ m).
Figure 9: Dependence of the torque acting on a particle from the liquid, relative to the center of the nanotube, on $x_{C}$ coordinate of the nanotube ($\alpha=\pi/3$, $y_{C}=0.5\cdot 10^{-3}$ m).
The rotation angle of the nanotube during particle motion in liquid, while the fixing radius has not been reached, can be determined by solving the system of differential equations,
@@
Fig. 10 shows the result of a numerical calculation of the change in angle $\alpha$ (in degrees) over time. The calculation was carried out until the particle reached the fixation radius. Calculations for different initial positions of the particle showed that the angle of rotation of the nanotube during its movement in the liquid does not exceed $5^{\circ}$ (this situation is shown in Fig. 10).
@@
Consider the rotation of a nanotube by the fluid flow when the particle touches the fixing radius with its edge, let this point be $A_{1}$. In this case, it ceases to move translationally and performs only rotational motion relative to point $A_{1}$. The rotational moment can be approximately determined based on the formula (12),
@@
where $I_{A1}=ml^{2}/3$ is the moment of inertia of the particle about the axis perpendicular to the nanotube and passing through its end $A_{1}$, $\vec{\omega}$ is the angular velocity of rotation, and $\vec{M}$ is the torque exerted on the particle by the liquid relative to point $A_{1}$. We approximate the torque considering that the nanotube is small and the averaged torque is applied to its center, point $C$,
@@
Integrating over time, assuming the vector $\overrightarrow{A_{1}C}$ is constant over the small time interval $\Delta t$, we obtain the approximate formula for the $i$-th time step,
@@
$\frac{ml^{2}}{3}\left.\omega_{z}\right|_{i}\approx\left.m\left(\left(x_{C}-x_{A_{1}}\right)v_{y}(x_{C},\,y_{C})-\left(y_{C}-y_{A_{1}}\right)v_{x}(x_{C},\,y_{C})\right)\right|_{i}.$ (21)
If the nanotube has touched another particle at its point $N$, then the rotational moment from the liquid relative to point $N$ is approximately given by (similarly to formula (17))
$\vec{M}=\int_{A_{1}N}\left[\overrightarrow{NA},\,dm\left.\frac{d\vec{v}}{dt}\right|_{A}\right]+\int_{NA_{2}}\left[\overrightarrow{NA},\,dm\left.\frac{d\vec{v}}{dt}\right|_{A}\right]\approx\\ \approx\left[\overrightarrow{NS_{1}},\,\left.\frac{d\vec{v}}{dt}\right|_{S_{1}}\right]\frac{m}{l}\cdot|NA_{1}|+\left[\overrightarrow{NS_{2}},\,\left.\frac{d\vec{v}}{dt}\right|_{S_{2}}\right]\frac{m}{l}\cdot|NA_{2}|,$ (22)
@@
The equation of the rotational motion of the nanotube relative to the point $N$ is
@@
where $I_{N}=ml^{2}/12+m\left|CN\right|^{2}$ is the moment of inertia of the particle relative to the axis be perpendicular to the nanotube and passing through the point $N$,
$I_{N}\left.\vec{\omega}\right|_{i}\approx\left(\left.\left[\overrightarrow{NS_{1}},\,\left.\vec{v}\right|_{S_{1}}\right]\frac{m}{l}\cdot|NA_{1}|+\left[\overrightarrow{NS_{2}},\,\left.\vec{v}\right|_{S_{2}}\right]\frac{m}{l}\cdot|NA_{2}|\right)\right|_{i},$
$\left(ml^{2}/12+m\left|CN\right|^{2}\right)\left.\omega_{z}\right|_{i}\approx\frac{m}{l}\left(\left|N\,A_{1}\right|\cdot\left(\left(x_{S_{1}}-x_{N}\right)v_{y}\left(x_{S_{1}},\,y_{S_{1}}\right)-\left(y_{S_{1}}-y_{N}\right)v_{x}(x_{S_{1}},\,y_{S_{1}})\right)\right.+\\ +\left.\left.\left|N\,A_{2}\right|\cdot\left(\left(x_{S_{2}}-x_{N}\right)v_{y}(x_{S_{2}},\,y_{S_{2}})-\left(y_{S_{2}}-y_{N}\right)v_{x}(x_{S_{2}},\,y_{S_{2}})\right)\right)\right|_{i},$ (24)
$\left|N\,A_{1}\right|=\sqrt{\left(x_{A1}-x_{N}\right)^{2}+\left(y_{A1}-y_{N}\right)^{2}},\qquad\left|N\,A_{2}\right|=\sqrt{\left(x_{A2}-x_{N}\right)^{2}+\left(y_{A2}-y_{N}\right)^{2}},$

### v2

@@ IV Conclusion
The results of the numerical simulations indicate that the combined effect of advection, diffusion, and electrostatics, under certain values of key system parameters ($N_{p}$, $t_{\mathrm{max}}$, and $kq^{2}$) leads to a transition from order to disorder in the morphology of the nanotube ring deposit, associated with the coffee-ring effect, which was previously observed in the experiment [8]. The concentration of particles affects the average distance between nanotubes in a solution. The strength of the electrostatic interaction between the nanotubes depends on the distance between them and the magnitude of their charge. This interaction can induce mutual nanotube alignment during the initial stages of the process. The evaporation time affects the velocity of capillary flow, which transfers particles and rotates them when they reach the periphery of a droplet, where the thickness of the liquid layer corresponds to the diameter of nanotubes. This boundary is known as the fixing radius, $R_{f}$ [20]. Each nanotube that crosses this boundary is deposited on the substrate (i.e., it comes into contact with the substrate and sticks to it due to adhesion). In the model, the boundary $R_{f}$ moves as the liquid evaporates until it collapses at the central point of the region. In reality, the situation may be more complicated due to the appearance of many local ruptures (dry spots) in an ultrathin film. However, this does not affect the results of the numerical simulations presented here, as such local ruptures, if they occur, appear only after the ring-shaped deposit has been formed. For further research, a modified method with restricted evaporation is also interesting, when a droplet is located between two substrates [27]. This experiment showed the possibility of obtaining aligned nanotube precipitation with an area of several square centimeters. In addition, it is necessary to develop 3D models for a more accurate description of these processes. In certain cases, additional effects must be taken into account, such as sedimentation, contact line motion, and hydrodynamic and capillary interactions between particles — including those in binary mixtures with varying shapes (e.g. spheres and ellipsoids [28]) or sizes.
@@ Appendix A Rotation of the nanotube by the fluid flow
Consider the rotation relative to the center of the nanotube mass at point $C$. The moment of the force exerted on a particle by a liquid relative to a given point is equal to
@@
$\frac{dv_{x}}{dt}=\frac{R^{2}x}{4\left(x^{2}+y^{2}\right)\left(t_{\mathrm{max}}-t\right)^{2}}\left(\frac{1}{\sqrt{1-\frac{x^{2}+y^{2}}{R^{2}}}}-1+\frac{x^{2}+y^{2}}{R^{2}}\right),$ (14)
$\frac{dv_{y}}{dt}=\frac{R^{2}y}{4\left(x^{2}+y^{2}\right)\left(t_{\mathrm{max}}-t\right)^{2}}\left(\frac{1}{\sqrt{1-\frac{x^{2}+y^{2}}{R^{2}}}}-1+\frac{x^{2}+y^{2}}{R^{2}}\right).$ (15)
@@
The position of the particle is determined by the coordinates of its center of mass $(x_{C},\,y_{C})$ and the angle $\alpha$ between the nanotube and the $OX$ axis. Let us make the following substitutions: $y_{A}=\tan(\alpha)\times x_{A}+b,\quad dl=dx_{A}/\cos(\alpha),\quad\text{where}\quad b=y_{C}-\tan(\alpha)\times x_{C}$, and compute the integral,
$\vec{M}=\frac{\vec{n}_{z}R^{2}m/l}{4\left(t_{\mathrm{max}}-t\right)^{2}}\int_{x_{A_{1}}}^{x_{A_{2}}}\frac{b(x_{A}-x_{C})}{\left(\frac{x_{A}}{\cos(\alpha)}+b\sin(\alpha)\right)^{2}+b^{2}\cos^{2}(\alpha)}\times\\ \times\left(\frac{1}{\sqrt{1-\frac{\left(x_{A}/\cos(\alpha)+b\sin(\alpha)\right)^{2}+b^{2}\cos^{2}(\alpha)}{R^{2}}}}-1\right.+\\ +\left.\frac{\left(\frac{x_{A}}{\cos(\alpha)}+b\sin(\alpha)\right)^{2}+b^{2}\cos^{2}(\alpha)}{R^{2}}\right)\frac{dx_{A}}{\cos(\alpha)}=\frac{\vec{n}_{z}R^{2}b\,m/l}{4\left(t_{\mathrm{max}}-t\right)^{2}}\times\\ \times\left[\cos(\alpha)\left(f_{1}(g_{A1})-g_{A1}+\ln(g_{A1})-f_{1}(g_{A2})+g_{A2}-\ln(g_{A2})\right)\right.-\\ -\frac{b\sin(\alpha)\cos(\alpha)+x_{C}}{R\sqrt{W}}\left\{\arctan\left(\frac{0.5(Wg_{A2}-2W+g_{A2})}{\sqrt{W\left(1-g_{A2}\right)\left(g_{A2}-W\right)}}\right)\right.+\\ +f_{2}(g_{A2})-2\arctan\left(\frac{\sqrt{g_{A2}-W}}{\sqrt{W}}\right)-\\ -\arctan\left(\frac{Wg_{A1}-2W+g_{A1}}{2\sqrt{W\left(1-g_{A1}\right)\left(g_{A1}-W\right)}}\right)-\\ -\left.\left.f_{2}(g_{A1})+2\arctan\left(\frac{\sqrt{g_{A1}-W}}{\sqrt{W}}\right)\right\}\right],$ (17)
@@
$g_{A1}=\frac{\left(x_{A1}/\cos(\alpha)+b\sin(\alpha)\right)^{2}+b^{2}\cos^{2}(\alpha)}{R^{2}},$
$g_{A2}=\frac{\left(x_{A2}/\cos(\alpha)+b\sin(\alpha)\right)^{2}+b^{2}\cos^{2}(\alpha)}{R^{2}},$
$W=\left(\frac{b\cos(\alpha)}{R}\right)^{2},$
$f_{1}(\xi)=2\operatorname{arctanh}\sqrt{1-\xi},$
$f_{2}(\xi)=2\sqrt{W}\sqrt{\xi-W},$
$\xi$ is $g_{A1}$ or $g_{A2}$.
@@
$\vec{M}=\int_{A_{1}C}\left[\overrightarrow{CA},\,dm\left.\frac{d\vec{v}}{dt}\right|_{A}\right]+\int_{CA_{2}}\left[\overrightarrow{CA},\,dm\left.\frac{d\vec{v}}{dt}\right|_{A}\right]\approx\\ \approx\left[\overrightarrow{CS_{1}},\,\left.\frac{d\vec{v}}{dt}\right|_{S_{1}}\right]\frac{m}{2}+\left[\overrightarrow{CS_{2}},\,\left.\frac{d\vec{v}}{dt}\right|_{S_{2}}\right]\frac{m}{2}=\\ =\vec{n}_{z}\frac{mR^{2}}{8\left(t_{\mathrm{max}}-t\right)^{2}}\left(f_{3}(x_{S1},y_{S1})+f_{3}(x_{S2},y_{S2})\right),$ (18)
where $S_{1}$ and $S_{2}$ are some points on the nanotube located on opposite sides of its center,
$\hfill f_{3}(x,y)=\hfill\\ =\frac{y\left(x-x_{C}\right)-x\left(y-y_{C}\right)}{x^{2}+y^{2}}\left(\frac{1}{\sqrt{1-\frac{x^{2}+y^{2}}{R^{2}}}}-1+\frac{x^{2}+y^{2}}{R^{2}}\right).$
A comparative analysis of the formulas for the torque acting on a particle from the liquid relative to the center of a nanotube is presented in Figs. 8 and 9 for the time moment $t=100$ s. Note that $t_{\mathrm{max}}=200$ s, $M_{1}$ is calculated by the formula (17), $M_{2},\,M_{3}$ are calculated by the formula (18). For points $S_{1}$ and $S_{2}$ in $M_{2}$, the midpoints of the nanotube halves (points $E$ and $G$) have been used; in $M_{3}$, two points at the ends of the nanotube ($A_{1}$ and $A_{2}$) have been used. The torque calculated using the approximate formula (18) exhibits fairly good accuracy and can be employed for calculations in place of the exact formula (17), which slows down the computation process.
Figure 8: Dependence of the torque acting on a particle from the liquid, relative to the center of the nanotube, on the angle $\alpha$ ($x_{C}=10^{-3}$ m, $y_{C}=0.5\times 10^{-3}$ m).
Figure 9: Dependence of the torque acting on a particle from the liquid, relative to the center of the nanotube, on $x_{C}$ coordinate of the nanotube ($\alpha=\pi/3$, $y_{C}=0.5\times 10^{-3}$ m).
The rotation angle of the nanotube during particle motion in liquid, while the fixing radius has not been reached, can be determined by solving the system of differential equations:
@@
Figure  10 shows the result of a numerical calculation of the change in angle $\alpha$ (in degrees) over time. The calculation was carried out until the particle reached the fixing radius. The calculation results for different initial positions of the particle show that the rotation of the nanotube begins near the edge of the droplet, where the flow velocity is high. Notably, the angle of rotation does not exceed $5^{\circ}$. One of these calculation results is shown in Fig. 10.
@@
Consider the rotation of a nanotube by the fluid flow when the particle touches the fixing radius with its edge, and let this point be $A_{1}$. In this case, it ceases to move translationally and performs only rotational motion relative to point $A_{1}$. The rotational moment can be approximately determined based on the formula (13),
@@
where $I_{A1}=ml^{2}/3$ is the moment of inertia of the particle about the axis perpendicular to the nanotube and passing through its end $A_{1}$, $\vec{\omega}$ is the angular velocity of rotation, and $\vec{M}$ is the torque exerted on the particle by the liquid relative to point $A_{1}$. We approximate the torque considering that the nanotube is small and the averaged torque is applied to its center, point $C$:
@@
Integrating over time, assuming the vector $\overrightarrow{A_{1}C}$ is constant over the small time interval $\Delta t$ (numerical calculations of the nanotube rotation angle in $\Delta t=10^{-4}$ s around the point of contact demonstrate values less than 0.02), we obtain the approximate formula for the $i$-th time step:
@@
$\frac{ml^{2}}{3}\left.\omega_{z}\right|_{i}\approx\\ \approx\left.m\left(\left(x_{C}-x_{A_{1}}\right)v_{y}(x_{C},\,y_{C})-\left(y_{C}-y_{A_{1}}\right)v_{x}(x_{C},\,y_{C})\right)\right|_{i}.$ (22)
If the nanotube has touched another particle at its point $N$, then the rotational moment from the liquid relative to point $N$ is approximately given by [similarly to formula (18)]
$\vec{M}=\int_{A_{1}N}\left[\overrightarrow{NA},\,dm\left.\frac{d\vec{v}}{dt}\right|_{A}\right]+\int_{NA_{2}}\left[\overrightarrow{NA},\,dm\left.\frac{d\vec{v}}{dt}\right|_{A}\right]\approx\\ \approx\left[\overrightarrow{NS_{1}},\,\left.\frac{d\vec{v}}{dt}\right|_{S_{1}}\right]\frac{m}{l}\times|NA_{1}|+\left[\overrightarrow{NS_{2}},\,\left.\frac{d\vec{v}}{dt}\right|_{S_{2}}\right]\frac{m}{l}\times|NA_{2}|,$ (23)
@@
The equation of the rotational motion of the nanotube relative to point $N$ is
@@
where $I_{N}=ml^{2}/12+m\left|CN\right|^{2}$ is the moment of inertia of the particle relative to the axis that is perpendicular to the nanotube and passing through point $N$,
$I_{N}\left.\vec{\omega}\right|_{i}\approx\left(\left.\left[\overrightarrow{NS_{1}},\,\left.\vec{v}\right|_{S_{1}}\right]\frac{m}{l}\times|NA_{1}|+\left[\overrightarrow{NS_{2}},\,\left.\vec{v}\right|_{S_{2}}\right]\frac{m}{l}\times|NA_{2}|\right)\right|_{i},$
$\left(m\,l^{2}/12+m\left|CN\right|^{2}\right)\left.\omega_{z}\right|_{i}\approx\\ \approx\frac{m}{l}\left[\left|N\,A_{1}\right|\times f_{4}(x_{S1},y_{S1})\right.+\left.\left.\left|N\,A_{2}\right|\times f_{4}(x_{S2},y_{S2})\right]\right|_{i},$ (25)
$f_{4}(x,y)=\left(x-x_{N}\right)\,v_{y}\left(x,\,y\right)-\left(y-y_{N}\right)\,v_{x}(x,\,y),$
$\left|N\,A_{1}\right|=\sqrt{\left(x_{A1}-x_{N}\right)^{2}+\left(y_{A1}-y_{N}\right)^{2}},$
$\left|N\,A_{2}\right|=\sqrt{\left(x_{A2}-x_{N}\right)^{2}+\left(y_{A2}-y_{N}\right)^{2}},$
@@
Formula (22) is obtained from formula (25) in the limit when $N\to A_{1}$.